OpenStudy (anonymous):

Find the LCD of 1/3 x 3/4 and 11/16. Then rewrite it as an equivalent fraction.

OpenStudy (anonymous):

LCD as in lowest common denominator? First of all, simplify the first fraction and tell me what you get.

OpenStudy (mayankdevnani):

@EmoLover3

OpenStudy (anonymous):

@mayankdevnani Please don't just give answers, it's not even right.

OpenStudy (anonymous):

I still don't understand :/ i'm not good at this

OpenStudy (mayankdevnani):

answer is \[\frac{11}{64}\]

OpenStudy (mayankdevnani):

my calculation is wrong...first time @Traxter

OpenStudy (anonymous):

thx :)

OpenStudy (mayankdevnani):

welcome.....

OpenStudy (anonymous):

No it's not @mayankdevnani. @EmoLover3, you want to find the lowest common denominator that the fractions can have. The fractions we have are: \[\frac{1}{3}\times\frac{3}{4}=\frac{1}{4}\] and \[\frac{11}{16}\] To do this, you need to find the lowest common multiple of the denominators. The denominators are 4 and 16, so we need to think, what is the lowest number which both 4 and 16 divide into? The answer is 16, since 4x4=16 and 1x16=16. So the LCD is 16. Is there anything I can explain better for you? Don't listen to @mayankdevnani , I have no idea what he's talking about.

OpenStudy (anonymous):

As an equivalent fraction, we need to get 1/4 as a fraction wth 16 as the denominator. So we need to multiply on top and bottom by 4 to get: \[\frac{4}{16}\]